OpenStudy (anonymous):

3 tan x - cot x - 5 = 0 Need to use the quadratic formula to solve. The answer should be in radians (from the unit circle). I am not sure how to start, since the equation isn't in ax^2 + bx + c format.

OpenStudy (anonymous):

write cotx as 1/tanx and simplify

OpenStudy (anonymous):

I'm not supposed to simplify, I'm supposed to solve, meaning I should just be left with, say, tan x = 0. If I do get that, then I'd have to find where tan x is equal to 0 on the unit circle. That would be my answer. There will be more than one answer. I'm not sure I'm explaining it clearly...

OpenStudy (anonymous):

The instructions say "Solve for 0 ≤ x ≤ 2π"

OpenStudy (anonymous):

3 tan x - cot x - 5 = 0 3 tanx -1/tanx -5 =0 3tan^2(x) -1 - 5tanx =0 let tanx = t 3t^2 - 5t -1 =0

OpenStudy (anonymous):